OpenStudy (anonymous):

How many solutions are there to the following system of equations? 6x + 12y = 11 4x + 8y = 3

OpenStudy (***[isuru]***):

first lets take 6x + 12y = 11 divide the whole equation by 6 x + 2y = 11/6 ----(1) then , lets take 4x + 8y = 3 divide the whole equation by 4 x + 2y = 3/4 ----- (2) if u (2) - (1) it will be something like this.. 0 = 11/6 - 3/4 all the unknown variables r gone !! which means ... no solutions.. got it :) ?